# Chow polynomials of totally nonnegative matrices and posets

### v1

@@ Abstract.
Huh-Stevens and Ferroni-Schröter independently conjectured that the Hilbert-Poincaré series of the Chow ring of any geometric lattice has only real zeros, and Ferroni, Matherne and the second author extended this conjecture to the Chow polynomial of any Cohen-Macaulay poset. In this paper we address these conjectures by providing new defining relations and properties of Chow functions of posets and matrices. These are used, in conjunction with new techniques on interlacing sequences of polynomials, to prove that the Chow polynomial of any lower triangular and totally nonnegative matrix with all diagonal entries equal to one is real-rooted. This in turn is used to prove the above conjectures for a large class of posets and matroids that includes projective and affine geometries, dual partition and Dowling lattices, perfect matroid designs and paving matroids.
A consequence of our theory is a vast generalization of a recent theorem of Athanasiadis on real-rootedness preserving properties of the Eulerian transformation, which was conjectured by the first author and Jochemko.
We study Chow polynomials of Toeplitz matrices in greater detail. These turn out to be related to a family of generalized Eulerian polynomials with coefficients in the ring of symmetric polynomials that have been studied frequently in the literature, by e.g. Stanley, Brenti, Stembridge and Shareshian-Wachs. Hence the theory developed in this paper serves as a framework for these polynomials, which are proved to be real-rooted by our methods (when suitably specialized).
@@ 1. Introduction
In their influential work [FY04], Feichtner and Yuzvinsky introduced the Chow ring of an atomistic lattice $L$, and gave a presentation for it via relations in terms of chains in the lattice. This ring is always graded and finite-dimensional, allowing one to study its Hilbert-Poincaré series as an invariant of the lattice. When $L$ is a geometric lattice, then the Chow ring satisfies the Kähler package [AHK18], which implies that the Hilbert series of the Chow ring of a matroid is always palindromic and unimodal. Moreover for specific classes of matroids, these Hilbert series coincide with well-known families of polynomials; for example when $L$ is a Boolean algebra, then one recovers Eulerian polynomials, and when $L$ is a corank-one uniform geometric lattice, then one recovers the derangement polynomials [HRS21]. For these reasons, these Hilbert series have garnered importance on their own and are now known as the Chow polynomials of matroids. A tightly connected construction is the one of augmented Chow ring of $L$ introduced in [BHM+22], whose Hilbert series is known as the augmented Chow polynomial.
Huh-Stevens [Ste21, Conjectures 4.1.3 and 4.3.3] and Ferroni-Schröter [FS24, Conjecture 8.18] conjectured that Chow polynomials and augmented Chow polynomials of matroids have only real zeros. This was recently verified in the uniform case in [BV25, Theorem 1.1] and [FMSV24, Theorem 1.10], respectively. Ferroni, Matherne and the second author extended the definition of Chow polynomials to arbitrary bounded posets using the framework of Kazhdan-Lusztig-Stanley theory [FMV24]. These are palindromic polynomials that are conjectured to have only real zeros for all Cohen-Macaulay posets and Bruhat intervals (with the appropriate $P$-kernels [FMV24, Conjecture 4.26 and 6.14]). In this larger context, Hoster and Stump [HS25, Theorem 1.1] proved the real-rootedness for boolean complexes with nonnegative $h$-vectors.
The aim of this paper is to develop a systematic approach to the study of zeros of Chow polynomials for a class of posets called $\operatorname{\mathrm{TN}}$-posets, which were introduced in [BL24, BL25]. This class includes truncations of affine and projective geometries, perfect matroid designs, dual partition lattices and dual Dowling lattices. To do so, we define and study Chow polynomials of totally nonnegative matrices ($\operatorname{\mathrm{TN}}$-matrices), i.e., matrices whose minors are all nonnegative. Our main result is Theorem 4.16, which states that the Chow polynomial of any lower triangular $\operatorname{\mathrm{TN}}$-matrix with all diagonal entries equal to one is real-rooted. A direct consequence of this is Theorem 6.1, which says that the Chow polynomial of any $\operatorname{\mathrm{TN}}$-poset is real-rooted. We also use our methods to prove that the Chow polynomial of any paving matroid is real-rooted, Corollary 7.2.
@@
In Section 3 we use the machinery developed in Section 2 to define and study (augmented) Chow polynomials of lower triangular matrices $R$ with all diagonal entries equal to one. We relate this construction to (augmented) Chow polynomials of weak-rank uniform posets, a class of highly structured posets. We define and study a notion of $\mathcal{I}_{n}$-interlacing sequences of polynomials, which refines the notion of interlacing sequences to the case when the polynomials in question interlace their reciprocals. The new defining relations for Chow functions obtained in Section 2 translate into recursions that preserve $\mathcal{I}_{n}$-interlacing sequences of polynomials when combined with resolvability of $\operatorname{\mathrm{TN}}$-matrices (Definition 4.1). This is used to prove interlacing preserving properties of the so called Chow-deranged map and the Chow-Eulerian transformation associated to $R$. These are vast generalizations of two results of the first author and Solus [BS21], and Athanasiadis [Ath23], respectively, on real-rootedness preserving properties of the deranged map and the Eulerian transformation. Indeed Theorems 4.14 and 4.15 generalize these results to any lower triangular $\operatorname{\mathrm{TN}}$-matrix with all diagonal entries equal to one. A direct consequence of this is Theorem 4.16 which states that the Chow polynomials associated to any such matrix are real-rooted.
@@
In Section 5 we provide an explicit interpretation of the coefficients of the $\gamma$-polynomial of Chow polynomials of matrices as sums of certain minors of the matrix (Corollary 5.2) and as enumerators of certain non-intersecting paths in a network, following the work by Gessel and Viennot [GV85]. In Sections 6 and 7 we prove Theorem 6.1 and Theorem 7.1, which assert the real-rootedness of (augmented) Chow polynomials of $\operatorname{\mathrm{TN}}$-posets and generalized paving posets, a class of posets that contains lattices of flats of paving matroids.
Finally, in Section 8, we discuss the case of Toeplitz matrices and the related class of boolean posets, and provide generating functions for the Chow polynomials of interest in Theorems 8.1 and 8.5. In Theorem 8.4, we prove that the Chow polynomials associated to Toeplitz matrices of Pólya frequency sequences are real-rooted. We show that the Chow polynomials associated to the Toeplitz matrices
$(e_{i-j}(\mathbf{x}))_{i,j=0}^{\infty}\quad\mbox{ and }\quad(h_{i-j}(\mathbf{x}))_{i,j=0}^{\infty}$
@@
coincide with the generalized Eulerian polynomials, with symmetric function coefficients, that have been studied frequently in the literature, by Stanley, Brenti, Stembridge, Shareshian and Wachs and others, see [SW10] and the references therein. Hence the theory developed in this paper serves as a framework for these polynomials, that, for example, enables us to reprove (Schur) $\gamma$-positivity results of Gessel and Shareshian and Wachs, as well as real-rootedness for the polynomials obtained when $\mathbf{x}$ and $\mathbf{y}$ are chosen to be suitable real and nonnegative vectors.
@@ 2. The incidence algebra and Chow functions
The following theorem gives an alternative characterization of $P$-kernels.
@@
Notice that $\kappa-t\delta$ evaluated at $t=1$ is equal to $g(1)^{-1}g(1)-\delta=0$, and hence t1t-1 divides $\kappa-t\delta$ and thus $\overline{\kappa}$ is an element of $I_{\rho}(P)$ as claimed. The Chow function associated to $\kappa$ (or $g$) is then defined as
$H=(-\overline{\kappa})^{-1},$
i.e., the negative inverse of the reduced $P$-kernel with respect to the convolution product in the incidence algebra.
@@
Moreover, $H$ is the Chow function associated to $g$, and the polynomials $d_{x,y}$ satisfy the recursion
@@
We call the unique function $d$ achieved by Theorem thm: def H and d the Chow-derangement function associated to $g$. If the poset is bounded, we call $H_{P}=H_{\hat{0},\hat{1}}$ the Chow polynomial of $P$ (with respect to $g$) and $d_{P}=d_{\hat{0},\hat{1}}$ the Chow-derangement polynomial of $P$.
@@
$d_{n}=\sum_{\sigma\in\mathfrak{D}_{n}}t^{\mathrm{exc}(\sigma)},$
which count derangements (fixed point free permutations) in the symmetric group $\mathfrak{S}_{n}$ by number of excedances, and the Eulerian polynomials
@@
where $\mathrm{exc}(\sigma)=|\{i:\sigma(i)>i\}|$, see [Bre90, Corollary 1]. Hence we deduce that $d_{x}$ and $H_{x}$ are the traditional derangement polynomials and Eulerian polynomials, respectively.
@@
$H_{x,y}=g_{x,y}+t\sum_{\begin{subarray}{c}x\leq z<y\\ \rho_{x,z}\geq 2\end{subarray}}H_{\tau[x,z]}g_{z,y},\quad\mbox{ for all }x\leq y.$
@@
$H_{x,y}=1+t\sum_{\begin{subarray}{c}x<z\leq y\\ \rho_{x,z}\geq 2\end{subarray}}H_{\tau([x,z])}.$ (2.7)
@@ 2.1. Augmented Chow functions
In [FMV24, Definition 3.13] the augmented Chow function was defined as
$G=\mathcal{I}(g)H.$
We now give an alternative equivalent definition.
@@
Moreover $G$ is the augmented Chow function associated to $g$, and the polynomials $A_{x,y}$ satisfy the recursion
@@ 3. Chow polynomial of matrices and weak-rank uniform posets
$d_{n}=t\mathcal{S}_{n-1}\left(\sum_{k=0}^{n-1}r_{n,k}d_{k}\right),\quad 1\leq n\leq N.$
@@ 3.1. Weak-rank uniform posets
If $\rho(x)=n$, we define $r_{n,k}=|\{z\leq x:\rho(z)=k\}|$, and write $R=R(P)=(r_{n,k})_{n,k\geq 0}$. Notice that $r_{n,0}=r_{n,n}=1$ for each $n$. If $P$ is graded and satisfies (3.4), then we say that $P$ is rank-uniform.
@@
Let $P$ be a weak-rank uniform poset, and let $R=R(P)=(r_{n,k})_{n,k=0}^{N}$ be the corresponding matrix, where $N\in\mathbb{N}\cup\{\infty\}$. If $n\leq N$, then $H_{n}[R]$ is equal to the characteristic Chow polynomial $H_{x}$, where $x$ is any element of $P$ of rank $n$.
@@ 4.1. Totally nonnegative and resolvable matrices
  2. (ii)
    There are linear diagonal operators $\alpha_{i}:\mathbb{R}[t]\to\mathbb{R}[t]$, $1\leq i\leq N$, such that α_i(t^k)= α_i,kt^k, where α_i,k 0 for all i,k, and R_n = (t+ α_1)(t+α_2) (t+α_n) 1.
  3. (iii)
@@
4.2. Interlacing zeros and $\mathcal{I}_{n}$-interlacing sequences
Suppose $p$ and $q$ are two real-rooted polynomials in $\mathbb{R}[t]$ with zeros $\cdots\leq\beta_{2}\leq\beta_{1}$ and $\cdots\leq\alpha_{2}\leq\alpha_{1}$, respectively. The zeros interlace if
@@ 4.2. Interlacing zeros and $\mathcal{I}_{n}$-interlacing sequences
We say that a sequence of polynomials $f_{1},f_{2},\ldots,f_{m}\in\mathbb{R}_{\geq 0}[t]$ of degree at most $n$ is $\mathcal{I}_{n}$-interlacing if
@@
Suppose $f_{1},f_{2},\ldots,f_{m}$ is an interlacing ($\mathcal{I}_{n}$-interlacing) sequence. Then the sequences
@@
We claim that either $\mathcal{S}_{n}(f)\equiv 0$, or the leading coefficient of $\mathcal{S}_{n}(f)$ is positive. Since $f\prec\mathcal{I}_{n}(f)$, the degree of $f$ is smaller or equal to the degree of $\mathcal{I}_{n}(f)$. If the degree of $\mathcal{I}_{n}(f)$ is strictly larger than the degree of $f$, then the leading coefficient of $\mathcal{S}_{n}(f)$ equals that of $\mathcal{I}_{n}(f)$ and is thus positive. Otherwise write $f=t^{M}h$, where $h(0)\neq 0$, and suppose $h$ has degree m=nMNm=n-M-N. Then
$\mathcal{I}_{n}(f)=t^{N}\mathcal{I}_{n-M-N}(h).$
Since the degrees of $f$ and $\mathcal{I}_{n}(f)$ are equal, it follows that $M=N$. Also $h\in\mathbb{I}_{m}$, and the leading coefficient of $\mathcal{S}_{n}(f)$ is equal to the leading coefficient of $\mathcal{S}_{m}(h)$. Write
@@
$\alpha_{m}\geq 1/\alpha_{1}\geq\alpha_{m-1}\geq 1/\alpha_{2}\geq\cdots\geq\alpha_{1}\geq 1/\alpha_{m}>0.$
Hence
@@
which follows from Lemmas 4.6 and 4.11. Hence $g_{0},\ldots,g_{m+1}$ is interlacing. Notice that for polynomials of degree at most $d$, $f\prec g$ if and only if $\mathcal{I}_{d}(g)\prec\mathcal{I}_{d}(f)$. Since $\mathcal{I}_{n+1}(g_{m+1})=g_{m+1}$,

### v2

@@ Abstract.
Huh-Stevens and Ferroni-Schröter independently conjectured that Hilbert-Poincaré series of Chow rings of geometric lattices have only real zeros. Ferroni, Matherne and the second author extended this conjecture to Chow polynomials of Cohen-Macaulay poset. In this paper we address the above conjectures by providing new defining relations and properties of Chow functions of posets and matrices. These are used, in conjunction with new techniques on interlacing sequences of polynomials, to prove that Chow polynomials of totally nonnegative matrices have only real zeros, which, in turn, proves the above conjectures for a class of posets that contains projective and affine geometries, face lattices of cubical polytopes, dual partition and Dowling lattices, perfect matroid designs, and lattices of flats of paving matroids.
We also study Chow polynomials of Toeplitz matrices in greater detail, and show how these are related the combinatorics of binomial and Sheffer posets, as well as to a family of generalized Eulerian polynomials with coefficients in the ring of symmetric polynomials that have been studied by e.g. Stanley, Brenti, Stembridge and Shareshian-Wachs.
Key words and phrases:
Totally nonnegative matrix, real-rooted polynomial, Pólya frequency sequence, Chow polynomial, Chow ring, geometric lattice, binomial poset, Sheffer poset
2020 Mathematics Subject Classification:
06A07, 15B48, 13D40, 26C10
PB is a Wallenberg Academy Scholar supported by the Knut and Alice Wallenberg Foundation, and the Göran Gustafsson foundation.
@@ 1. Introduction
In their influential work [FY04], Feichtner and Yuzvinsky introduced the Chow ring of an atomistic lattice $L$, and gave a presentation for it via relations in terms of chains in the lattice. This ring is graded and finite-dimensional, allowing one to study its Hilbert-Poincaré series as an invariant of the lattice. When $L$ is a geometric lattice, then the Chow ring satisfies the Kähler package [AHK18], which implies that the Hilbert series of the Chow ring of a geometric lattice is always palindromic and unimodal. Moreover for specific classes of geometric lattices, these Hilbert series coincide with well-known families of polynomials; for example for boolean algebras one recovers the Eulerian polynomials, and for truncations of boolean algebras one recovers the derangement polynomials [HRS21]. For these reasons, the Hilbert series of Chow rings of geometric lattices (or matroids) have garnered importance on their own and are now known as the Chow polynomials of matroids. A tightly connected construction is the one of augmented Chow ring of $L$ introduced in [BHM+22], whose Hilbert series is known as the augmented Chow polynomial.
Ferroni-Schröter [FS24, Conjecture 8.18] and Huh-Stevens [Ste21, Conjectures 4.1.3 and 4.3.3] independently conjectured that Chow polynomials and augmented Chow polynomials of matroids have only real zeros. This was recently verified in the uniform case in [BV25, Theorem 1.1] and [FMSV24, Theorem 1.10], respectively. Ferroni, Matherne and the second author extended the definition of Chow polynomials to arbitrary bounded posets using the framework of Kazhdan-Lusztig-Stanley theory [FMV24]. These are palindromic polynomials that are conjectured to have only real zeros for all Cohen-Macaulay posets and Bruhat intervals (with the appropriate $P$-kernels [FMV24, Conjectures 4.26 and 6.14]). In this larger context, Hoster and Stump [HS25, Theorem 1.1] proved that Chow polynomials of boolean complexes with nonnegative $h$-vectors are real-rooted.
In this paper we develop a systematic approach to the study of zeros of Chow polynomials for a class of posets called $\operatorname{\mathrm{TN}}$-posets, which were introduced in [BL24, BL25]. This class includes (rank-selected subposets of) affine and projective geometries, perfect matroid designs, dual partition lattices and dual Dowling lattices. To do so, we define and study Chow polynomials of totally nonnegative matrices ($\operatorname{\mathrm{TN}}$-matrices), i.e., matrices whose minors are all nonnegative. Our main result is Theorem 4.18, which states that the Chow polynomial of any lower triangular $\operatorname{\mathrm{TN}}$-matrix with all diagonal entries equal to one is real-rooted. A direct consequence of this is Theorem 6.1, which says that the Chow polynomial of any $\operatorname{\mathrm{TN}}$-poset is real-rooted. We also use our methods to prove that the Chow polynomial of any paving matroid is real-rooted, Corollary 7.2.
@@
In Section 3 we use the machinery developed in Section 2 to define and study (augmented) Chow polynomials of matrices. We relate this construction to (augmented) Chow polynomials of weak-rank uniform posets, a class of highly structured posets. We define and study a notion of $\mathcal{I}_{n}$-interlacing sequences of polynomials, which refines the notion of interlacing sequences to the case when the zeros of the polynomial in question interlace the zeros of its reciprocal. The new defining relations for Chow functions obtained in Section 2 translate into recursions that preserve $\mathcal{I}_{n}$-interlacing sequences of polynomials when combined with resolvability of $\operatorname{\mathrm{TN}}$-matrices (Definition 4.1), thus refining the method using resolvability that was initiated in [BL24]. This is used to prove interlacing preserving properties of the so called Chow-deranged map and the Chow-Eulerian transformation associated to a matrix $R$. These are vast generalizations of two results of the first author and Solus [BS21], and Athanasiadis [Ath23], respectively, on real-rootedness preserving properties of the deranged map and the Eulerian transformation. Indeed Theorems 4.16 and 4.17 generalize these results to any lower triangular $\operatorname{\mathrm{TN}}$-matrix with all diagonal entries equal to one. A direct consequence is Theorem 4.18 which states that the Chow polynomials associated to any such matrix are real-rooted.
@@
In Section 5 we provide an explicit interpretation of the coefficients of the $\gamma$-polynomial of Chow polynomials of matrices as sums of certain minors of the matrix (Corollary 5.2) and as enumerators of certain non-intersecting paths in a network, following the work by Gessel and Viennot [GV85]. In Sections 6 and 7 we prove Theorem 6.1 and Theorem 7.1, which assert that (augmented) Chow polynomials of $\operatorname{\mathrm{TN}}$-posets and generalized paving posets (a class of posets that contains lattices of flats of paving matroids) are real-rooted.
In Section 8, we study the case of Toeplitz matrices in greater detail, and provide generating function identities for the Chow polynomials of interest in Theorem 8.1. In Theorem 8.4, we prove that the Chow polynomials associated to Toeplitz matrices of Pólya frequency sequences are real-rooted. In Section 8.1 we provide expressions for the generating series for the various Chow polynomials in the case when $P$ is a binomial poset or a Sheffer poset. These are classes of highly regular posets studied by Doubilet, Rota and Stanley [DRS72], and Ehrenborg and Readdy [ER95], respectively.
@@
In Section 8.2 we show that the Chow polynomials associated to the Toeplitz matrices $(e_{i-j}(\mathbf{x}))_{i,j=0}^{\infty}$ and $(h_{i-j}(\mathbf{x}))_{i,j=0}^{\infty}$ coincide with the generalized Eulerian polynomials, with symmetric function coefficients, that have been studied frequently in the literature, by Stanley, Brenti, Stembridge, Shareshian and Wachs and others, see [SW10] and the references therein. Hence the theory developed in this paper serves as a framework for these polynomials, that, for example, enables us to reprove Schur $\gamma$-positivity results of Gessel and Shareshian-Wachs, as well as real-rootedness for the polynomials obtained when $\mathbf{x}$ and $\mathbf{y}$ are chosen to be suitable real and nonnegative vectors.
@@ 2. The incidence algebra and Chow functions
The following theorem gives a characterization of $P$-kernels.
@@
Notice that $\kappa-t\delta$ evaluated at $t=1$ is equal to $g(1)^{-1}g(1)-\delta=0$, and hence t1t-1 divides $\kappa-t\delta$ and thus $\overline{\kappa}$ is an element of $I_{\rho}(P)$ as claimed. The Chow function associated to $\kappa$ (or $g$) is defined as the element of $I_{\rho}(P)$,
$(-\overline{\kappa})^{-1},$ (2.2)
i.e., the negative inverse of the reduced $P$-kernel.
@@
Moreover, $H$ is the Chow function (2.2) associated to $g$, and the polynomials $d_{x,y}$ satisfy the recursion
@@
We call the unique function $d$ achieved by Theorem 2.2 the Chow-derangement function associated to $g$. If the poset is bounded, we call $H_{P}=H_{\hat{0},\hat{1}}$ the Chow polynomial of $P$ (with respect to $g$) and $d_{P}=d_{\hat{0},\hat{1}}$ the Chow-derangement polynomial of $P$.
@@
$d_{n}=\sum_{\sigma\in\mathfrak{D}_{n}}t^{\mathrm{exc}(\sigma)},\quad\quad\mathrm{exc}(\sigma)=|\{i:\sigma(i)>i\}|,$
which count derangements (fixed point free permutations) in the symmetric group $\mathfrak{S}_{n}$ by the number of excedances, and the Eulerian polynomials
@@
see [Bre90, Corollary 1]. Hence we deduce that $d_{x}$ and $H_{x}$ are the traditional derangement polynomials and Eulerian polynomials, respectively.
@@
$H_{x,y}=g_{x,y}+t\sum_{\begin{subarray}{c}x\leq z\leq y\\ \rho_{x,z}\geq 2\end{subarray}}H_{\tau[x,z]}g_{z,y},\quad\mbox{ for all }x\leq y.$
@@
$H_{x,y}=1+t\sum_{\begin{subarray}{c}x\leq z\leq y\\ \rho_{x,z}\geq 2\end{subarray}}H_{\tau([x,z])}.$ (2.8)
@@ 2.1. Augmented Chow functions
In [FMV24, Definition 3.13] the augmented Chow function was defined as the element in $I_{\rho}(P)$,
$\mathcal{I}(g)H.$ (2.9)
The next theorem offers an alternative equivalent definition.
@@
Moreover $G$ is the augmented Chow function (2.9) associated to $g$, and the polynomials $A_{x,y}$ satisfy the recursion
@@ 3. Chow polynomial of matrices and weak-rank uniform posets
$d_{n}=t\mathcal{S}_{n-1}\left(\sum_{k=0}^{n-1}r_{n,k}d_{k}\right),\quad 1\leq n\leq N.$ (3.1)
@@ 3.1. Weak-rank uniform posets
If $\rho(x)=n$, we define $r_{n,k}=r_{n,k}(P)=|\{z\leq x:\rho(z)=k\}|$, and write $R=R(P)=(r_{n,k})_{n,k\geq 0}$. Notice that $r_{n,0}=r_{n,n}=1$ for each $n$. If $P$ is graded and satisfies (3.5), then we say that $P$ is rank-uniform.
@@
Let $P$ be a weak-rank uniform poset, and let $R=(r_{n,k}(P))_{n,k=0}^{N}$ be the corresponding matrix, where $N\in\mathbb{N}\cup\{\infty\}$. If $n\leq N$, then $H_{n}[R]$ is equal to the characteristic Chow polynomial $H_{x}$, where $x$ is any element of $P$ of rank $n$.
@@ 4.1. Totally nonnegative and resolvable matrices
  • (ii)
    There are linear diagonal operators $\alpha_{i}:\mathbb{R}[t]\to\mathbb{R}[t]$, $1\leq i\leq N$, such that
    $\alpha_{i}(t^{k})=\alpha_{i,k}t^{k},\mbox{ where }\alpha_{i,k}\geq 0\mbox{ for all }i,k,$
    and
    $R_{n}=(t+\alpha_{1})(t+\alpha_{2})\cdots(t+\alpha_{n})1.$
  • (iii)
@@
4.2. Interlacing zeros and $\mathcal{I}_{n}$-interlacing sequences of polynomials
Recall that a polynomial $p\in\mathbb{R}[t]$ is called real-rooted if all of its zeros are real. For technical reasons, the zero polynomial is considered to be real-rooted. Suppose $p$ and $q$ are two real-rooted polynomials in $\mathbb{R}[t]$ with zeros $\cdots\leq\beta_{2}\leq\beta_{1}$ and $\cdots\leq\alpha_{2}\leq\alpha_{1}$, respectively. The zeros interlace if
@@ 4.2. Interlacing zeros and $\mathcal{I}_{n}$-interlacing sequences of polynomials
We will now define a central notion of this paper.
Definition 4.9.
We say that a sequence $f_{1},f_{2},\ldots,f_{m}$ of polynomials in $\mathbb{R}_{\geq 0}[t]$ of degree at most $n$ is $\mathcal{I}_{n}$-interlacing if
@@
Suppose $f_{1},f_{2},\ldots,f_{m}$ is an interlacing ($\mathcal{I}_{n}$-interlacing) sequence of polynomials. Then the sequences
@@
Suppose $f,g$ is an $\mathcal{I}_{n}$-interlacing sequence. We claim that either $\mathcal{S}_{n}(f)\equiv 0$, or the leading coefficient of $\mathcal{S}_{n}(f)$ is positive. Since $f\prec\mathcal{I}_{n}(f)$, the degree of $f$ is smaller or equal to the degree of $\mathcal{I}_{n}(f)$. If the degree of $\mathcal{I}_{n}(f)$ is strictly smaller than the degree of $f$, then the leading coefficient of $\mathcal{S}_{n}(f)$ equals that of $\mathcal{I}_{n}(f)$ and is thus positive.
Otherwise if the degrees of $f$ and $\mathcal{I}_{n}(f)$ are equal, then write $f=t^{M}h$ where $h(0)\neq 0$, and suppose $h$ has degree m=nMNm=n-M-N. Then
$\mathcal{I}_{n}(f)=t^{N}\mathcal{I}_{n-M-N}(h)=t^{N}\mathcal{I}_{m}(h).$
Since the degrees of $f$ and $\mathcal{I}_{n}(f)$ agree, it follows that $M=N$. Also $h\in\mathbb{I}_{m}$, and since $\mathcal{S}_{n}(f)=t^{N}\mathcal{S}_{m}(h)$, the leading coefficient of $\mathcal{S}_{n}(f)$ is equal to the leading coefficient of $\mathcal{S}_{m}(h)$. Write
@@
$\alpha_{m}\geq 1/\alpha_{1}\geq\alpha_{m-1}\geq 1/\alpha_{2}\geq\cdots\geq\alpha_{1}\geq 1/\alpha_{m}>0,$
since $h\in\mathbb{I}_{m}$. From this we deduce
@@
which follows from Lemmas 4.6 and 4.12. Hence $g_{0},\ldots,g_{m+1}$ is interlacing.
Notice that for polynomials of degree at most $d$, $f\prec g$ if and only if $\mathcal{I}_{d}(g)\prec\mathcal{I}_{d}(f)$. Since $\mathcal{I}_{n+1}(g_{m+1})=g_{m+1}$,